# RLCAD: Reinforcement Learning Training Gym for Revolution Involved CAD Command Sequence Generation

### v2

@@ 1 Introduction
Over the past few years, generative models have flourished, demonstrating immense potential. Representative approaches include diffusion models [1], GANs [2], autoencoders [3], large language models (LLMs) [4, 5], and vision-language models (VLMs) [6]. These methods have been extensively applied in generating CAD models, such as boundary representation (B-Rep) autoregressive encoder BrepGen [7], end-to-end point cloud-to-command sequence generation model CADSIGNet [8], text-to-command sequence generation model Text2CAD [9], multi-modal command sequence generation model CAD-MLLM [10], B-Rep to command sequence generation model Fusion 360 Gallery [11], voxel-to-command sequence generation model SECAD-Net [12] and SfmCAD [13], autoregressive command sequence generation models HNC-CAD [14], SkexGen [15], and DeepCAD [16], as well as LLM based CAD code generation algorithm [17] and platform Zoo.dev [18]. Despite their success, most of these methods support only basic operations like 2D sketch, extrusion, and Boolean operations.
@@
Traditional CAD geometry generation algorithms rely on CAD geometric engines, which provide essential functionalities such as validity checks and constraint solving – capabilities that supervised learning alone cannot achieve. Meanwhile, SOTA large language models such as DeepSeek R1 [19] and OpenAI o1 [20] employ reinforcement learning (RL) techniques to enhance text generation quality and model reasoning capabilities. In bionic robotics, RL-based simulation training has also yielded impressive results in robot motion control [21] and manipulation [22].
Inspired by these advancements, we present RLCAD, a CAD RL training gym based on CAD geometric engine. In this framework, the policy network of the RL framework generates a command sequence, which is executed in the gym to produce the corresponding CAD geometry. This geometry is then fed back into the policy network to calculate rewards and update the network.
@@
Based on this training gym, we propose an RL-based CAD geometry generation algorithm that converts B-Rep models into command sequences. Our approach supports operations beyond sketches, Boolean, and extrusion, including revolution operations. As shown in Figure 1, including revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, chess pieces, etc. We summarize our contributions as follows:
  • 1.
    We present a CAD RL training gym bridging RL algorithms with CAD geometric engines. We provide a high-level Python interface wrapper and support parallel environment sampling, enabling efficient interaction with multiple CAD environments. In addition, we integrate the RL algorithm library Stable Baseline3 [23] and the OpenAI Gymnasium interface [24] with our training gym, making it easier for developers to implement and test RL algorithms.
  • 2.
@@
    We present an RL-based CAD model generation algorithm that converts B-Rep models into command sequences within our training gym. Our approach features a multi-modal policy network that integrates multiple common CAD operations with cross-modal feature extractors. To further enhance generation quality, we introduce a set of hybrid reward functions that jointly guide the learning process.
@@ 2.1 CAD Datasets
Current parametric CAD models typically are stored as B-Rep [25] or command sequence [26]. B-Rep includes the analytical representation of the surface and curves of a 3D shape, while command sequences record the historical construction process of the shape. B-Rep models consist of trimmed parametric surfaces along with topological information describing their adjacency relationships.
The ABC dataset [27] is the most widely used B-Rep dataset in academia, containing one million CAD models. DeepCAD [16] cleaned 170,000 pairs of B-Rep and command sequences from the ABC dataset. Text2CAD [9] constructed hierarchical descriptions for the DeepCAD dataset, ranging from coarse to fine granularity, using LLMs and VLMs, adding text descriptions to the command sequences. CAD-MLLM [10], using a method similar to DeepCAD, extracted 185,000 models from the ABC dataset and generated corresponding text descriptions, multi-view images, and point clouds for these models. However, these datasets primarily contain 2D sketches, such as lines, arcs, and circles, along with extrusion and Boolean operations.
@@ 2.2 CAD Model Generation
BrepGen [7] combines vector variational autoencoders and latent space diffusion models for the autoregressive generation of B-Rep models. CADSIGNet [8] employs layer-wise cross-attention between point clouds and command sequences to learn implicit representations of CAD models, enabling end-to-end point cloud to command sequence generation. HNC-CAD [14], SkexGen [15], and DeepCAD [16] quantize and encode CAD command sequences, generating corresponding latent space features and training through an encoder-decoder autoregressive model. SfmCAD [13] uses voxels as input and proposes differentiable sketch and sweeping path modeling operations to reconstruct the CAD model in a self-supervised manner.
@@ 2.3 B-Rep to Command Sequence Generation
Several commercial CAD softwares [28] employ heuristic feature extraction to generate command sequences from B-Rep models. However, the 2D sketches extracted using this method are often structurally complex and difficult to edit further. Fusion 360 Gallery [11] is a representative deep learning-based approach that first extracts features to obtain continuous geometric parameters for basic operations (e.g., 2D sketching, extrusion, etc.). It then uses imitation learning to generate confidence scores of the command sequence from the input B-Rep model. Finally, a confidence-based local search algorithm, combined with geometric similarity comparison, optimizes model accuracy in a simulation environment. However, this method only supports face-extrusion operations, and the local search algorithm heavily relies on the accuracy of the confidence scores from supervised learning.
@@ 2.4 RL Gym
RL involves updating states and receiving rewards or penalties through agent-environment interactions to maximize long-term returns, thereby learning decision-making strategies [29]. Based on environmental modeling methods, RL can be divided into model-based and model-free categories.
Model-based approaches first model the state transition function and reward function of the environment using small-scale datasets, then optimize the policy network under the RL framework using the environment model [30]. A representative method is RL with Human Feedback [31], widely used in pre-training large language models. Model-free methods replace the environment model with real-world environments or physical simulators, obtaining immediate feedback on the policy network through real-time interaction within the RL framework [32]. These methods are widely applied in bionic robot training simulations, such as robot motion control and dexterous hand grasping [21, 22]. Representative robot simulation engines include Isaac [33], MuJoCo [34], and Genesis [35].
Therefore, we propose to build a CAD training gym based on the CAD geometric engine, Parasolid [36], which can be easily extended to integrate with other CAD engines, such as OpenCascade [37] or ACIS [38].
Figure 2: The training pipeline is composed of two stages. In the first stage, a contrastive learning approach is employed to pre-train the UV-Net network, aiming to derive an encoder model that can effectively characterize the B-Rep of CAD model. During the second stage, a reinforcement learning approach is employed to generate the command sequence. We first utilize the tunable UV-Net model to extract the B-Rep embedding of the CAD model, which is then integrated with the feature vector of the historical modeling action sequence. Subsequently, the Actor-Critic network predicts the action distribution and value. The predicted action is transmitted to RLCADGym for execution, yielding the next-stage observation. The neural reward and geometric reward are designed to update the policy network.
@@ 3.1 Our Approach
The overall training pipeline is depicted in Figure 2. In the first stage, we adopt a contrastive learning approach to pre-train UV-Net [39] on a large-scale dataset of unlabeled CAD models, aiming to derive an embedding that can effectively characterize the high-dimensional B-Rep feature of CAD model. Given a B-Rep model represented as a face adjacency graph, UV-Net processes it through a combination of surface convolution and curve convolution operations, followed by a Graph Neural Network (GNN) to extract a compact and discriminative embedding. To enhance the robustness of the learned representations, we apply UV-Net’s data augmentation strategies, including connected patch extraction and stochastic node/edge dropping.
In the second training stage, the whole network can be divided into two parts: the feature extraction network and the Actor-Critic network. Inspired by advancements in natural language processing (NLP), we consider incorporating the action sequence as part of the state and fusing it with B-Rep features using cross-attention. This allows us to measure the gap between the current action sequence and the target geometry. Specifically, for the current B-Rep and target B-Rep, we use the pre-trained UV-Net to extract features from the face adjacency graphs, obtaining $g_{c}$ and $g_{t}$. These features are stacked to form $G_{\text{stack}}$, and self-attention is applied to fuse $g_{c}$ and $g_{t}$, resulting in $\tilde{G}_{\text{stack}}$. For the action sequence, we use an $L$-layer GTrXL network to extract features. And then we take the last valid action $h_{\text{last\_valid}}$ as the feature of the action sequence. Cross-attention is then applied to fuse $h_{\text{last\_valid}}$ and $\tilde{G}_{\text{stack}}$. Finally, the fused feature is served as input to the Actor and Critic networks, predicting the probability distribution of actions and the cumulative reward of the current state, respectively. The action is sampled from the probability distribution and interacts with RLCADGym to obtain the reward and the next state. Through multiple rounds of interaction and network parameters update, the optimal action sequence is predicted.
@@ 3.2 Preliminary
We formulate the CAD reconstruction task from B-Rep to command sequence as a Markov Decision Process (MDP). The key components of this formulation are defined as follows: States $(s)\in\mathcal{S}$ contains the current and target B-Rep geometry, we use a face-adjacency graph [39] to represent the B-Rep geometry. Action $(a)\in\mathcal{A}$ represents a modeling operation that allows the simulator to modify the current geometry with extrusion, revolution, and Boolean operations. State transition kernel $T$ allows the simulator to apply the modeling operation to update the current geometry. Reward $(r)\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ allows the simulator to calculate the difference between the current and target states, and reward discount $\gamma\in[0,1)$, determining the importance of future rewards.
@@ 4.1 Face-Extrusion Operation
[…]
@@ 5 Generation of CAD Command Sequence
In this section, we will delve into the training process of command sequence generation. Extensive experiments across different RL frameworks indicate that the Proximal Policy Optimization (PPO) algorithm [40] produces the highest-quality command sequences.
@@ 5.1 Proximal Policy Optimization
where $\widehat{A}_{t}$ is computed via Generalized Advantage Estimation (GAE), which combines the advantages of Monte Carlo estimation and Temporal Difference (TD) methods, effectively balancing TD error and variance [41]:
@@ 5.2 Face Adjacency Graph
We systematically convert B-Rep models into attributed graphs through three sequential stages: surface feature extraction, adjacency construction, and graph normalization. Let $G=(\mathcal{V},\mathcal{E})$ denotes the graph where nodes $v_{i}\in\mathcal{V}$ represent CAD faces and edges $e_{ij}\in\mathcal{E}$ encode face adjacency [39].
@@
Edges are derived from B-Rep topology links to encode global structure. For each directional link $l\in\mathcal{L}$, bidirectional edges are created by validating node existence: =$\mathcal{E}=\left\{(i,j)\,|\,\exists\,l:v_{i}=\text{source}(l),\,v_{j}=\text{target}(l),\,v_{i},v_{j}\in\mathcal{V}\right\}\cup\left\{(j,i)\,|\,(i,j)\in\mathcal{E}\right\}$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$ is symmetrized through $\mathbf{A}=\max(\mathbf{A}_{\text{raw}},\mathbf{A}_{\text{raw}}^{\top})$, where $\mathbf{A}_{\text{raw}}$ represents the initial directional adjacency.
Graph normalization stabilizes training by augmenting self-loops and applying row-wise scaling. Self-connections are added via the identity matrix $\mathbf{I}$, and degree normalization is performed using the degree matrix $\mathbf{D}$, where $D_{ii}=\sum_{j}(A_{ij}+\delta_{ij})$ with $\delta_{ij}$ as the Kronecker delta [42]. The normalized adjacency is computed as $\widehat{\mathbf{A}}=\mathbf{D}^{-1}(\mathbf{A}+\mathbf{I})$, ensuring $\forall i,\sum_{j}\widehat{A}_{ij}=1\,$. Isolated nodes are handled by zeroing invalid degrees, and the result is stored in sparse coordinate format for computational efficiency.

### v3

@@ 1 Introduction
Over the past few years, generative models have flourished, demonstrating immense potential. Representative approaches include diffusion models [dhariwal2021diffusion], GANs [mirza2014conditional], autoencoders [kingma2013auto], large language models (LLMs) [vaswani2017attention, devlin2018bert], and vision-language models (VLMs) [chen2022pali]. These methods have been extensively applied in generating CAD models, such as boundary representation (B-Rep) autoregressive encoder BrepGen [xu2024brepgen], end-to-end point cloud-to-command sequence generation model CADSIGNet [khan2024cad], text-to-command sequence generation model Text2CAD [khan2024text2cad], multi-modal command sequence generation model CAD-MLLM [xu2024cad], B-Rep to command sequence generation model Fusion 360 Gallery [willis2021fusion], voxel-to-command sequence generation model SECAD-Net [li2023secad] and SfmCAD [li2024sfmcad], autoregressive command sequence generation models HNC-CAD [xu2023hierarchical], SkexGen [xu2022skexgen], and DeepCAD [wu2021deepcad], as well as LLM based CAD code generation algorithm [alrashedy2024generating] and platform Zoo.dev [zoo_dev]. Despite their success, most of these methods support only basic operations like 2D sketch, extrusion, and Boolean operations.
@@
Traditional CAD geometry generation algorithms rely on CAD geometric engines, which provide essential functionalities such as validity checks and constraint solving – capabilities that supervised learning alone cannot achieve. Meanwhile, SOTA large language models such as DeepSeek R1 [guo2025deepseek] and OpenAI o1 [jaech2024openai] employ reinforcement learning (RL) techniques to enhance text generation quality and model reasoning capabilities. In bionic robotics, RL-based simulation training has also yielded impressive results in robot motion control [zhao2024zsl] and manipulation [fu2024mobile].
Inspired by these advancements, we present the first RL training gym tailored to parametric CAD command sequence generation on exact B-Rep geometry. We provide a high-level Python interface wrapper and support parallel environment sampling, enabling efficient interaction with multiple CAD environments. In this framework, first we represent the B-Rep as a face-adjacency graph, then employ a graph neural network based feature extractor to encode the B-Rep, and finally use a policy network to predict an action. Each action is described by four elements—the start face, end face, operation type (newbody, intersection, union, or subtraction), and action type (extrude or revolve), where the start and end faces differ for extrusion but coincide for revolution—which is executed in the gym to produce the corresponding CAD geometry. The resulting geometry is then fed back into the policy network to compute rewards and update the network parameters.
@@
Our approach supports operations beyond sketches, Boolean, and extrusion, including revolution operations. As shown in Figure 1, including revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, chess pieces, etc.
Technical contributions
@@
We present a CAD-focused RL training gym and an RL-driven CAD model generation algorithm that converts B-Rep models into command sequences within the training gym. Our approach features a multi-modal policy network that integrates multiple common CAD operations with cross-modal feature extractors. To further enhance generation quality, we introduce a set of hybrid reward functions, including geometric similarity and neural similarity, which jointly guide the learning process and help better distinguish models in both geometric and semantic aspects.
We construct an RL-oriented dataset based on the ABC dataset [koch2019abc], applying filtering and de-duplication to obtain a curated set of $20k$ CAD models that contain both extrusion and revolution operations, together with their ground-truth command sequences. Unlike existing extrusion-dominated CAD benchmarks, our dataset expands the operational repertoire and is intended to support future research on general CAD command-sequence generation.
@@ 2.1 CAD Datasets
Current parametric CAD models typically are stored as B-Rep [braid1975synthesis] or command sequence [price1982design]. B-Rep includes the analytical representation of the surface and curves of a 3D shape, while command sequences record the historical construction process of the shape. B-Rep models consist of trimmed parametric surfaces along with topological information describing their adjacency relationships.
The ABC dataset [koch2019abc] is the most widely used B-Rep dataset in academia, containing one million CAD models. DeepCAD [wu2021deepcad] cleaned 170,000 pairs of B-Rep and command sequences from the ABC dataset. Text2CAD [khan2024text2cad] constructed hierarchical descriptions for the DeepCAD dataset, ranging from coarse to fine granularity, using LLMs and VLMs, adding text descriptions to the command sequences. CAD-MLLM [xu2024cad], using a method similar to DeepCAD, extracted 185,000 models from the ABC dataset and generated corresponding text descriptions, multi-view images, and point clouds for these models. However, these datasets primarily contain 2D sketches, such as lines, arcs, and circles, along with extrusion and Boolean operations.
@@ 2.2 CAD Model Generation
BrepGen [xu2024brepgen] combines vector variational autoencoders and latent space diffusion models for the autoregressive generation of B-Rep models. CADSIGNet [khan2024cad] employs layer-wise cross-attention between point clouds and command sequences to learn implicit representations of CAD models, enabling end-to-end point cloud to command sequence generation. HNC-CAD [xu2023hierarchical], SkexGen [xu2022skexgen], and DeepCAD [wu2021deepcad] quantize and encode CAD command sequences, generating corresponding latent space features and training through an encoder-decoder autoregressive model. SfmCAD [li2024sfmcad] uses voxels as input and proposes differentiable sketch and sweeping path modeling operations to reconstruct the CAD model in a self-supervised manner.
@@ 2.3 B-Rep to Command Sequence Generation
Several commercial CAD softwares [solidworks, inventor] employ heuristic feature extraction to generate command sequences from B-Rep models. However, the 2D sketches extracted using this method are often structurally complex and difficult to edit further. Fusion 360 Gallery [willis2021fusion] is a representative deep learning-based approach that first extracts features to obtain continuous geometric parameters for basic operations (e.g., 2D sketching, extrusion, etc.). It then uses imitation learning to generate confidence scores of the command sequence from the input B-Rep model. Finally, a confidence-based local search algorithm, combined with geometric similarity comparison, optimizes model accuracy in a simulation environment. However, this method only supports face-extrusion operations, and the local search algorithm heavily relies on the accuracy of the confidence scores from supervised learning.
@@ 2.4 RL Gym
RL involves updating states and receiving rewards or penalties through agent-environment interactions to maximize long-term returns, thereby learning decision-making strategies [wiering2012reinforcement]. Based on environmental modeling methods, RL can be divided into model-based and model-free categories.
Model-based approaches first model the state transition function and reward function of the environment using small-scale datasets, then optimize the policy network under the RL framework using the environment model [moerland2023model]. A representative method is RL with Human Feedback [christiano2017deep], widely used in pre-training large language models. Model-free methods replace the environment model with real-world environments or physical simulators, obtaining immediate feedback on the policy network through real-time interaction within the RL framework [ccalicsir2019model]. These methods are widely applied in bionic robot training simulations, such as robot motion control and dexterous hand grasping [zhao2024zsl, fu2024mobile]. Representative robot simulation engines include Isaac [makoviychuk2021isaac], MuJoCo [todorov2012mujoco], and Genesis [genesis].
Therefore, we propose to build a CAD training gym based on the CAD geometric engine, Parasolid [ParasolidWikipedia], which can be easily extended to integrate with other CAD engines, such as OpenCascade [opencascade] or ACIS [acis].
Figure 2: The training pipeline is composed of two stages. In the first stage, a contrastive learning approach is employed to pre-train the UV-Net network, aiming to derive an encoder model that can effectively characterize the B-Rep of the CAD model. During the second stage, a reinforcement learning approach is employed to generate the command sequence. We first utilize the tunable UV-Net model to extract the B-Rep embedding of the CAD model, which is then integrated with the feature vector of the historical modeling action sequence. Subsequently, the Actor-Critic network predicts the action distribution and value. The predicted action is transmitted to RLCADGym for execution, yielding the next-stage observation. The neural reward and geometric reward are designed to update the policy network.
@@ 3.1 Our Approach
The overall training pipeline is depicted in Figure 2. In the first stage, we adopt a contrastive learning approach to pre-train UV-Net [jayaraman2021uv] on a large-scale dataset of unlabeled CAD models, aiming to derive an embedding that can effectively characterize the high-dimensional B-Rep feature of CAD models. Given a B-Rep model represented as a face adjacency graph, UV-Net processes it through a combination of surface convolution and curve convolution operations, followed by a Graph Neural Network (GNN) to extract a compact and discriminative embedding. To enhance the robustness of the learned representations, we apply UV-Net’s data augmentation strategies, including connected patch extraction and stochastic node/edge dropping.
In the second training stage, the whole network can be divided into two parts: the feature extraction network and the Actor-Critic network. Inspired by advancements in natural language processing (NLP), we consider incorporating the action sequence as part of the state and fusing it with B-Rep features using cross-attention. This allows us to measure the gap between the current action sequence and the target geometry. Specifically, for the current B-Rep and target B-Rep, we use the pre-trained UV-Net to extract features from the face adjacency graphs, obtaining $g_{c}$ and $g_{t}$. These features are stacked to form $G_{\text{stack}}$, and self-attention is applied to fuse $g_{c}$ and $g_{t}$, resulting in $\tilde{G}_{\text{stack}}$. For the action sequence, we use an $L$-layer GTrXL network to extract features. And then we take the last valid action $h_{\text{last\_valid}}$ as the feature of the action sequence. Cross-attention is then applied to fuse $h_{\text{last\_valid}}$ and $\tilde{G}_{\text{stack}}$. Finally, the fused feature serves as input to the Actor and Critic networks, predicting the probability distribution of actions and the cumulative reward of the current state, respectively. The action is sampled from the probability distribution and interacts with RLCADGym to obtain the reward and the next state. Through multiple rounds of interaction and network parameters update, the optimal action sequence is predicted.
@@ 3.2 Preliminary
We formulate the CAD reconstruction task from B-Rep to a command sequence as a Markov Decision Process (MDP). The key components of this formulation are defined as follows: States $(s)\in\mathcal{S}$ contains the current and target B-Rep geometry, we use a face-adjacency graph [jayaraman2021uv] to represent the B-Rep geometry. Action $(a)\in\mathcal{A}$ represents a modeling operation that allows the simulator to modify the current geometry with extrusion, revolution, and Boolean operations. State transition kernel $T$ allows the simulator to apply the modeling operation to update the current geometry. Reward $(r)\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ allows the simulator to calculate the difference between the current and target states, and reward discount $\gamma\in[0,1)$, determining the importance of future rewards.
@@ 4.1 Face-Extrusion Operation
As shown in Figure 3, the extrusion operation takes a sketch as the initial geometry and then extrudes it from 2D to 3D. In the extrusion process, we need to specify a distance parameter to determine how far the profile is extruded along the normal direction. Additionally, a Boolean operator is required to specify whether the operation is newbody, intersection, union, or subtraction. Thus, our operation can be represented as $\{\text{start face},\text{end face},\text{op}\}$. Here, the start face and end face are a set of parallel planes on the target object, and op represents the type of Boolean operation. The start face defines the sketch for the extrusion operation, while the end face marks the extrusion distance. Therefore, the shape of the end face does not need to be identical to the start face; it only needs to be parallel to it.
@@ 4.2 Revolution Operation
The revolution operation takes a profile as the revolution surface, a line coplanar with the profile as the axis of rotation, and a rotation angle, then rotates the 2D profile around the axis under the rotation angle to form a 3D shape. We select a curved surface on the target geometry as the resulting surface of the revolve operation. The operation can be represented as $\{\text{face},\text{op}\}$, where face represents the revolution surface, and op denotes the type of Boolean operation.
@@ 5 Generation of CAD Command Sequence
In this section, we will delve into the training process of command sequence generation. Extensive experiments across different RL frameworks indicate that the Proximal Policy Optimization (PPO) algorithm [schulman2017proximal] produces the highest-quality command sequences.
@@ 5.1 Proximal Policy Optimization
where $\widehat{A}_{t}$ is computed via Generalized Advantage Estimation (GAE), which combines the advantages of Monte Carlo estimation and Temporal Difference (TD) methods, effectively balancing TD error and variance [SMLJA16]:
@@ 5.2 Face Adjacency Graph
We systematically convert B-Rep models into attributed graphs through three sequential stages: surface feature extraction, adjacency construction, and graph normalization. Let $G=(\mathcal{V},\mathcal{E})$ denotes the graph where nodes $v_{i}\in\mathcal{V}$ represent CAD faces and edges $e_{ij}\in\mathcal{E}$ encode face adjacency [jayaraman2021uv].
@@
Edges are derived from B-Rep topology links to encode global structure. For each directional link $l\in\mathcal{L}$, bidirectional edges are created by validating node existence: $\mathcal{E}=\left\{(i,j)\,|\,\exists\,l:v_{i}=\text{source}(l),\,v_{j}=\text{target}(l),\,v_{i},v_{j}\in\mathcal{V}\right\}\cup\left\{(j,i)\,|\,(i,j)\in\mathcal{E}\right\}$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$ is symmetrized through $\mathbf{A}=\max(\mathbf{A}_{\text{raw}},\mathbf{A}_{\text{raw}}^{\top})$, where $\mathbf{A}_{\text{raw}}$ represents the initial directional adjacency.
Graph normalization stabilizes training by augmenting self-loops and applying row-wise scaling. Self-connections are added via the identity matrix $\mathbf{I}$, and degree normalization is performed using the degree matrix $\mathbf{D}$, where $D_{ii}=\sum_{j}(A_{ij}+\delta_{ij})$ with $\delta_{ij}$ as the Kronecker delta [kozen2007indefinite]. The normalized adjacency is computed as $\widehat{\mathbf{A}}=\mathbf{D}^{-1}(\mathbf{A}+\mathbf{I})$, ensuring $\forall i,\sum_{j}\widehat{A}_{ij}=1\,$. Isolated nodes are handled by zeroing invalid degrees, and the result is stored in sparse coordinate format for computational efficiency.